OpenStudy (anonymous):

Solve the expression Ra^t = Qg^t for t assuming that all other letters are positive constants.

OpenStudy (anonymous):

R\Q=(g\a)^t log(R\Q)=tlog(g\a) t=log(r\q)\log(g\a)

OpenStudy (anonymous):

@mjbmanny Thanks!

OpenStudy (anonymous):

thanks for the medal.